# AEANet: Affinity Enhanced Attentional Networks for Arbitrary Style Transfer

### v1

@@ Abstract
Arbitrary artistic style transfer is a field that integrates rational academic research with emotional artistic creation. It aims to produce an image that not only features artistic characteristics of the target style but also preserves the texture structure of the content image itself. Existing style transfer methods primarily rely either on global statistics-based information or local patch-based. As a result, the generated images often either superficially apply a filter to the content image or capture extraneous semantic information from the style image, leading to a significant deviation from the global style. In this paper, we propose Affinity Enhanced-Attentional Networks (AEANet), which include a content affinity-enhanced attention (CAEA) module, style affinity-enhanced attention (SAEA) module, and hybrid attention (HA) module. The CAEA and SAEA modules first use attention to improve content and style representations with a Detail Enhanced(DE) module to reinforce fine details. Then, it aligns the global statistical information of the content and style features to fine-tune the feature information. Subsequently, the HA module adjusts the distribution of style features based on the distribution of content features. Additionally, we introduce affinity attention-based Local Dissimilarity Loss to preserve the affinities between the content and style images. Experimental results demonstrate that our approach outperforms state-of-the-art methods
Keywords: arbitrary style transfer, attention, local statistics swap, detail enhanced
@@ 1 Introduction
Scientific research is rational and rigorous, artistic creation is emotional and dynamic, and artistic style migration is a study that can perfectly balance scientific research and artistic creation. This study aims to disassemble the color and texture features in a stylized image and migrate the features to a given content image to achieve image style conversion while preserving the features of the content image
@@
in arbitrary style transfer. as much as possible under the constraints of style migration. In recent years, Gatys et al. [9] first used pre-trained deep convolutional neural networks [26] to separate and recombine images’ content and style features. They innovatively used the Gram matrix to quantify the correlations of style features, capturing and transferring the color and texture layout of the target style image. However, the method proposed by Gatys et al. [9] relies on an iterative gradient descent process to optimize the image, resulting in prolonged image generation speeds.
To expand the application scenarios of style transfer, some researchers have used feedforward network-based methods[16, 18, 2, 19, 27, 31, 24] to accomplish this task efficiently. However, these methods can only achieve the transfer of a specific style. Another group of researchers has been striving to extend the applicability of this research to arbitrary style transfer.[17, 20, 13, 1, 12, 6, 15] They focus on aligning global statistics by adjusting the overall distribution of the content image to match the second-order statistics of the style image. Huang et al. [13] adjusted the feature statistics (mean and variance) of the content image to match the statistical characteristics of the target style image. Li et al. [20] used whitening and coloring transformations to ensure that the feature covariance of the content image directly matches the statistical characteristics of the target style image. Specifically, the WCT [20] method removes the variance influence of content features through a whitening step and then transforms these features into a space that matches the covariance of the style image through a coloring process. This process goes beyond simple variance matching, achieving deeper style embedding. However, the high computational cost of WCT [20] limits its efficiency and real-time application.
Correspondingly, local patch-based methods [5, 10, 22, 25, 3, 8, 30, 32, 14, 7] identify and extract corresponding feature regions in content and style images through local feature matching, employing refined fusion strategies such as local linear transformation or feature map-based region matching. Park et al. [22] introduced Style-Attentional Networks, which use self-attention mechanisms to capture the correlation between style and content images. They effectively and flexibly decorate local style patterns based on the semantic spatial distribution of the content image. Although SANet achieves a good balance between style and content to some extent through its Identity Loss Function and feature embedding, it is challenging to find the optimal style transfer effect or preserve content details during the style transfer process for images with rich content details.
@@
  • We design the CAEA module, SAEA module, and HA module to capture local feature information, and to match global statistical information, and to enhance the detail representation in both content and style images, improving the effect of image stylization significantly.
@@ 2.1 Global Statistics-based Arbitrary Style Transfer
Global statistics-based methods typically achieve style transfer by adjusting the statistical properties of content features to align with the global statistical characteristics of style features. Li et al. [20] employed Whitening and Coloring Transform (WCT) to map content image features into a space where the covariance matches the style image’s. Huang et al. [13] adapted the mean and variance of content features using an adaptive instance normalization(AdaIN) layer to match the mean and variance of style features globally. Jing et al. [15] extended Huang’s work by introducing a Dynamic Instance Normalization (DIN) module, which uses learnable convolutions to encode the style image and stylize the content image accordingly. An et al. [1] proposed an unbiased style transfer framework that addressed the content leakage issue in existing general style transfer methods (e.g., WCT [20] or AdaIN [13]) by achieving unbiased image style transfer through reversible neural flows [12]. Chen et al. [4] first establish a global style by aligning the global statistics of content features and style features (mean and variance) and then enrich local style details by exchanging local statistical data. These methods have successfully achieved global distribution alignment between stylized images and the overall style of their corresponding style images. However, they overlook a critical issue: a style image generally encompasses multiple
Figure 2: Overview of our framework. We input the content image $I_{c}$ and the style image $I_{s}$ into a pre-trained VGG encoder, generating the corresponding features $F_{c}$ and $F_{s}$. These features $F_{c}$ and $F_{s}$ are then fed into the CAEA and SAEA modules, respectively. The HA module adjusts the style features $F_{ss}$ based on the distribution of the content features $F_{cc}$ to obtain the stylized features $F_{cs}$, which are decoded by a decoder symmetrically designed to the encoder to generate the resulting image $I_{cs}$.
stylistic regions, while a content image often contains several distinct semantic areas. Global style transfer needs to fully account for the diversity of local styles and the complexity of multiple semantic features in the content image.
@@ 3 Method
Firstly, we briefly introduce the workflow of our model in Section 3.1. Secondly, we explain our proposed Affinity Enhanced Attentional Network (AEANet) module in Section 3.2. Finally, we describe the Loss functions used in AEANet in Section 3.3, including our proposed Local Dissimilarity Loss based on affinity-enhanced attention.
@@ 3.1 Overview of AEANet
Generally, our work involves taking any given style image $I_{s}$ and content image $I_{c}$ and aiming to develop a generative model capable of synthesizing a stylized image $F_{cs}$. This model preserves the texture structure of the target image $I_{c}$ and learns local and global style features from the style image $I_{s}$. To achieve arbitrary artistic style transfer, we propose the Affinity Enhanced Attentional Network (AEANet). The AEANet is composed of a Content Affinity Enhanced Attentional (CAEA) module, Style Affinity Enhanced Attentional (SAEA) module, and Hybrid Attention (HA) module. As shown in Figure 2, the encoder $E$ and decoder $D$ are symmetrically designed based on the VGG-19 network [26], the network inputs are the content image $I_{c}$ and the style image $I_{s}$.
@@
3.2 Affinity Enhanced Attentional Network
The core concept of our AEANet involves passing the feature maps $(F_{c},F_{s})$, extracted by a pre-trained VGG-19 network, into the Content Affinity-Enhanced Attentional module and Style Affinity Enhanced Attentional module, respectively. Through two independent CAEA and SAEA modules, we perform localized feature extraction and detail enhancement for style and content, gaining $F_{cc}$ and $F_{ss}$. These outputs are subsequently fed into Hybrid Attention module, where they are recombined to produce $F_{cs}$.
3.2.1 Content Affinity Enhanced Attentional
The Content Affinity Enhanced Attentional module is divided into two steps: (1) calculating an attention map by utilizing deep and shallow features of the content image; (2) enhancing the details of the content image.
Calculate Attention Map Generally, Softmax attention can be expressed as.
@@ 3.2.1 Content Affinity Enhanced Attentional
Content Detail Enhanced We begin by subtracting the square of the attention score matrix $F_{\text{aff}}$ from itself, passing the result through the ReLU activation function, and then taking the square root of this matrix; This process yields a new matrix—weight matrix $M_{\text{weight}}$, which maintains the original matrix’s shape and has values distributed between 0 and 1. Subsequently, we standardize $F_{\text{cv}}$, multiply it by the detailed weight matrix $M_{\text{weight}}$, and finally add it to the original $F_{\text{aff}}$ matrix.The specific formula is as follows:
@@ 3.3 Loss Function
3.3.1 Perceptual Loss
Similar to AdaIN [13], we minimize the loss $L_{c}$ between the content image $I_{c}$ and the stylized image $I_{cs}$, as well as the loss $L_{s}$ between the style image $I_{s}$ and $I_{cs}$.
@@ 3.3.1 Perceptual Loss
Identity Loss follow SANet [22], We introduce identity loss to better preserve both content and style features.
@@
Table 1: Quantitative Evaluation. Comparison of style transfer models based on $L_{c}$ and $L_{s}$ performance metrics, we evaluate the effectiveness of various methods in preserving input content and style by calculating the average content and style loss values. SSIM [29] is a metric used to measure the similarity between two images, in style transfer, a higher value indicates a higher degree of similarity, we calculate the average SSIM using the results of stylization generated from the same content image and 60 different style photos. We use $512\times 512$ pixel images to calculate the average generation time. The $\uparrow$ indicates that a higher value signifies better performance, whereas a $\downarrow$ indicates that a lower value is better. The best results are underlined, and the second best is in bold.
@@ 4.2 Comparison with State-of-the-Art Methods
Qualitative Evaluation:Figure 3 presents a visual comparison of our results with eight different state-of-the-art (SOTA) models. It can be observed that AdaIN [13] captures less information from the style image due to the simplified alignment of mean and variance, resulting in the presence of colors in the generated image that are not found in either the content or style images (e.g.,$4^{th}$,$5^{th}$ and $6^{th}$rows). SANet [22] excessively introduces semantic structures from the style image (e.g.,$3^{th}$,$5^{th}$,$7^{th}$ and $9^{th}$ rows). Although MAST [8] employs an attention mechanism, its application does not yield the expected results. Analyses indicate that while the main structure of the content image remains clear, other structural parts are compromised. The style features in the generated images are insufficiently represented (e.g.,$1^{th}$,$2^{th}$,$5^{th}$ and $8^{th}$ rows). For TSSAT [4], despite balancing global statistics and local features, the preservation of the content image structure is suboptimal, resulting in some unappealing outcomes (e.g.,$1^{th}$,$3^{th}$,$8^{th}$ and $9^{th}$ rows). IEST [3] extracts fewer features from the content image, which leads to subpar preservation of the content structure in the stylized images (e.g.,$1^{th}$,$3^{th}$ and $4^{th}$ rows). The StyTr2 [7] method replaces the traditional VGG [26] encoder with a Transformer [28] encoder, which mitigates the drawbacks of reduced receptive fields during the encoding phase. However, content image information loss is still evident (e.g.,$1^{th}$,$2^{th}$ and $8^{th}$ rows). The ArtFlow [1] the model produces stylized images during inference that exhibit noticeable deviations from the input style images (e.g.,$1^{th}$,$3^{th}$,$7^{th}$ and $9^{th}$ rows). Consequently, our research findings demonstrate an effective preservation of content structure integrity and achieve desirable expression of style features.
@@ 5 Conclusion
In this paper, we propose AEANet to achieve high-quality style transfer. Initially, the model performs a local style feature information exchange and detail enhancement on both the content and style images. This is followed by a global statistical alignment to fine-tune the features. The HA module adjusts the style features based on the distribution of the fine-tuned content features. Our model achieves a good balance between preserving the structural texture of the content image and stylization. Experimental results demonstrate that our network fully considers local style features and incorporates global statistical information. The generated results attest to the superiority of our model. The main limitation of our study is the room for improvement in the output speed of high-resolution images. In future work, we aim to simplify computational complexity to enhance output efficiency.

### v2

# AEANet: Affinity-Enhanced Attentional Networks for Arbitrary Style Transfer

@@ Abstract
Arbitrary artistic style transfer is a research area that combines rational academic study with emotive artistic creation. It aims to create a new image from a content image according to a target artistic style, maintaining the content’s textural structural information while incorporating the artistic characteristics of the style image. However, existing style transfer methods often significantly damage the texture lines of the content image during the style transformation. To address these issues, we propose affinity-enhanced attentional network, which include the content affinity-enhanced attention (CAEA) module, the style affinity-enhanced attention (SAEA) module, and the hybrid attention (HA) module. The CAEA and SAEA modules first use attention to enhance content and style representations, followed by a detail enhanced (DE) module to reinforce detail features. The hybrid attention module adjusts the style feature distribution based on the content feature distribution. We also introduce the local dissimilarity loss based on affinity attention, which better preserves the affinity with content and style images. Experiments demonstrate that our work achieves better results in arbitrary style transfer than other state-of-the-art methods.
Keywords: arbitrary style transfer, attention, local patch-based, detail enhanced
@@ 1 Introduction
Scientific research is rational and rigorous, artistic creation is emotional and dynamic, and artistic style migration is a study that can perfectly balance scientific research and artistic creation. This study aims to disassemble the color and texture
@@
features in a stylized image and migrate the features to a given content image to achieve image style conversion while preserving the features of the content image as much as possible under the constraints of style migration. In recent years, Gatys et al. [9] first used pre-trained deep convolutional neural networks [26] to separate and recombine images’ content and style features. They innovatively used the Gram matrix to quantify the correlations of style features, capturing and transferring the color and texture layout of the target style image. However, the method proposed by Gatys et al. [9] relies on an iterative gradient descent process to optimize the image, resulting in prolonged image generation speeds.
To expand the application scenarios of style transfer, some researchers have used feedforward network-based methods [16, 18, 2, 19, 27, 31, 24] to accomplish this task efficiently. However, these methods can only achieve the transfer of a specific style. Another group of researchers has been striving to extend the applicability of this research to arbitrary style transfer. [17, 20, 13, 1, 12, 6, 15] They focus on aligning global statistics by adjusting the overall distribution of the content image to match the second-order statistics of the style image. Huang et al. [13] adjusted the feature statistics (mean and variance) of the content image to match the statistical characteristics of the target style image. Li et al. [20] used whitening and coloring transforms to ensure that the feature covariance of the content image directly matches the statistical characteristics of the target style image. Specifically, the WCT [20] method removes the variance influence of content features through a whitening step and then transforms these features into a space that matches the covariance of the style image through a coloring process. This process goes beyond simple variance matching, achieving deeper style embedding. However, the high computational cost of WCT [20] limits its efficiency and real-time application. Correspondingly, local patch-based methods [5, 10, 22, 25, 3, 8, 30, 32, 14, 7] identify and extract corresponding feature regions in content and style images through local feature matching, employing refined fusion strategies such as local linear transformation or feature map-based region matching. Park et al. [22] introduced style-attentional network (SANet) , which use self-attention mechanisms to capture the correlation between style and content images. They effectively and flexibly decorate local style patterns based on the semantic spatial distribution of the content image. Although SANet achieves a good balance between style and content to some extent through its identity loss function and feature embedding, it is challenging to find the optimal style transfer effect or preserve content details during the style transfer process for images with rich content details.
@@
  • We design the CAEA module, SAEA module and HA module to capture local feature information, and to match global statistical information, and to enhance the detail representation in both content and style images, improving the effect of image stylization significantly.
@@ 2.1 Global statistics-based arbitrary style transfer
Global statistics-based methods typically achieve style transfer by adjusting the statistical properties of content features to align with the global statistical characteristics of style features. Li et al. [20] employed Whitening and Coloring Transform (WCT) to map content image features into a space where the covariance matches the style image’s. Huang et al. [13] adapted the mean and variance of content features using an adaptive instance normalization(AdaIN) layer to match the mean and variance of style features globally. Jing et al. [15] extended Huang’s work by introducing a Dynamic Instance Normalization (DIN) module, which uses learnable convolutions to encode the style image and stylize the content image accordingly. An et al. [1] proposed an unbiased style transfer framework that addressed the content leakage issue in existing general style transfer methods (e.g., WCT [20] or AdaIN [13]) by achieving unbiased image style transfer through reversible neural flows [12]. Chen et al. [4] first establish a global style by aligning the global statistics of content features and style features (mean and variance) and then enrich local style
Figure 2: Overview of our framework. We input the content image $I_{c}$ and the style image $I_{s}$ into a pre-trained VGG encoder, generating the corresponding features $F_{c}$ and $F_{s}$. These features $F_{c}$ and $F_{s}$ are fed into the CAEA and SAEA modules, respectively. The HA module adjusts the style features $F_{ss}$ based on the distribution of the content features $F_{cc}$ to obtain the stylized features $F_{cs}$, which are decoded by a decoder symmetrically designed to the encoder to generate the resulting image $I_{cs}$.
details by exchanging local statistical data. These methods have successfully achieved global distribution alignment between stylized images and the overall style of their corresponding style images. However, they overlook a critical issue: a style image generally encompasses multiple stylistic regions, while a content image often contains several distinct semantic areas. Global style transfer needs to fully account for the diversity of local styles and the complexity of multiple semantic features in the content image.
@@ 3 Method
Firstly, we briefly introduce the workflow of our model in Section 3.1. Secondly, we explain our proposed affinity-enhanced attentional network (AEANet) module in Section 3.2. Finally, we describe the loss functions used in AEANet in Section 3.3, including our proposed local dissimilarity loss based on affinity-enhanced attention.
@@ 3.1 Overview of AEANet
Generally, our work involves taking any given style image $I_{s}$ and content image $I_{c}$ and aiming to develop a generative model capable of synthesizing a stylized image $F_{cs}$. This model preserves the texture structure of the target image $I_{c}$ and learns local and global style features from the style image $I_{s}$. To achieve arbitrary artistic style transfer, we propose the affinity-enhanced attentional network (AEANet). The AEANet is composed of a content affinity-enhanced attentional (CAEA) module, style affinity-enhanced attentional (SAEA) module, and hybrid attention (HA) module. As shown in Figure 2, the encoder $E$ and decoder $D$ are symmetrically designed based on the VGG-19 network [26], the network inputs are the content image $I_{c}$ and the style image $I_{s}$.
@@
3.2 Affinity-enhanced attentional network
The core concept of our AEANet involves passing the feature maps $(F_{c},F_{s})$, extracted by a pre-trained VGG-19 network, into the content affinity-enhanced attentional module and style affinity-enhanced attentional module, respectively. Through two independent CAEA and SAEA modules, we perform localized feature extraction and detail enhanced for style and content, gaining $F_{cc}$ and $F_{ss}$. These outputs are subsequently fed into HA module, where they are recombined to produce $F_{cs}$.
3.2.1 Content affinity-enhanced attentional
The content affinity-enhanced attentional module is divided into two steps: (1) calculating an attention map by utilizing deep and shallow features of the content image; (2) enhancing the details of the content image.
Calculate attention map: Generally, Softmax attention can be expressed as.
@@ 3.2.1 Content affinity-enhanced attentional
Content detail-enhanced: We begin by subtracting the square of the attention score matrix $F_{\text{aff}}$ from itself, passing the result through the ReLU activation function, and then taking the square root of this matrix; This process yields a new matrix—weight matrix $M_{\text{weight}}$, which maintains the original matrix’s shape and has values distributed between 0 and 1. Subsequently, we standardize $F_{\text{cv}}$, multiply it by the detailed weight matrix $M_{\text{weight}}$, and finally add it to the original $F_{\text{aff}}$ matrix.The specific formula is as follows:
@@ 3.3 Loss function
Perceptual loss: Similar to AdaIN [13], we minimize the loss $L_{c}$ between the content image $I_{c}$ and the stylized image $I_{cs}$, as well as the loss $L_{s}$ between the style image $I_{s}$ and $I_{cs}$.
@@
Identity loss: follow SANet [22], We introduce identity loss to better preserve both content and style features.
@@
Table 1: Quantitative evaluation. Comparison of style transfer models based on $L_{c}$ and $L_{s}$ performance metrics, we evaluate the effectiveness of various methods in preserving input content and style by calculating the average content and style loss values. SSIM [29] is a metric used to measure the similarity between two images, in style transfer, a higher value indicates a higher degree of similarity, we calculate the average SSIM using the results of stylization generated from the same content image and 60 different style photos. We use $512\times 512$ pixel images to calculate the average generation time, we use the average time to generate 300 stylized photos. The $\uparrow$ indicates that a higher value signifies better performance, whereas a $\downarrow$ indicates that a lower value is better. The best results are underlined, and the second best is in bold.
@@ 4.2 Comparison with State-of-the-Art methods
Qualitative evaluation: Figure 3 presents a visual comparison of our results with eight different SOTA models. It can be observed that AdaIN [13] captures less information from the style image due to the simplified alignment of mean and variance, resulting in the presence of colors in the generated image that are not found in either the content or style images (e.g.,$4^{th}$,$5^{th}$ and $6^{th}$rows). SANet [22] excessively introduces semantic structures from the style image (e.g.,$3^{th}$,$5^{th}$,$7^{th}$ and $9^{th}$ rows). Although MAST [8] employs an attention mechanism, its application does not yield the expected results. Analyses indicate that while the main structure of the content image remains clear, other structural parts are compromised. The style features in the generated images are insufficiently represented (e.g.,$1^{th}$,$2^{th}$,$5^{th}$ and $8^{th}$ rows). For TSSAT [4], despite balancing global statistics and local features, the preservation of the content image structure is suboptimal, resulting in some unappealing outcomes (e.g.,$1^{th}$,$3^{th}$,$8^{th}$ and $9^{th}$ rows). IEST [3] extracts fewer features from the content image, which leads to subpar preservation of the content structure in the stylized images (e.g.,$1^{th}$,$3^{th}$ and $4^{th}$ rows). The StyTr2 [7] method replaces the traditional VGG [26] encoder with a Transformer [28] encoder, which mitigates the drawbacks of reduced receptive fields during the encoding phase. However, content image information loss is still evident (e.g.,$1^{th}$,$2^{th}$ and $8^{th}$ rows). The ArtFlow [1] the model produces stylized images during inference that exhibit noticeable deviations from the input style images (e.g.,$1^{th}$,$3^{th}$,$7^{th}$ and $9^{th}$ rows). Consequently, our research findings demonstrate an effective preservation of content structure integrity and achieve desirable expression of style features.
@@ 5 Conclusion
In this paper, we propose AEANet to achieve high-quality style transfer. Initially, the model performs a local style feature information exchange and detail enhanced on both the content and style images. The HA module adjusts the style features based on the distribution of the fine-tuned content features. Our model achieves a good balance between preserving the structural texture of the content image and stylization. Experimental results demonstrate that our network fully considers local content and style features. The generated results attest to the superiority of our model. The main limitation of our study is the room for improvement in the output speed of high-resolution images. In future work, we aim to simplify computational complexity to enhance output efficiency.